OpenStudy (anonymous):

how do I factor 81c^2-16

OpenStudy (anonymous):

what number squared is 16?

OpenStudy (anonymous):

4 right

OpenStudy (anonymous):

right and what would you square to get \(81c^2\) ?

OpenStudy (anonymous):

umm 9c^2

OpenStudy (anonymous):

well actually \(9c\)

OpenStudy (anonymous):

since \((9c)^2=81c^2\)

OpenStudy (anonymous):

now factor the difference of two squares as \[a^2-b^2=(a+b)(a-b)\] with \(a=9c\) and \(b=4\)

OpenStudy (anonymous):

is it 65 = (9+4) (9-4) right

OpenStudy (anonymous):

there is no 65 in the question, and you are missing a c

OpenStudy (anonymous):

\[81c^2-16=(9c+4)(9c-4)\]

OpenStudy (anonymous):

what do I do now I am lost

OpenStudy (anonymous):

hello there

OpenStudy (anonymous):

help me please